OpenStudy (anonymous):

A painter leans a 25-foot ladder against the wall. The bottom of the ladder is 20 feet from the base of the wall. The painter then moves the bottom of the ladder 13 feet closer to the wall. How much higher is the top of the ladder on the wall?

OpenStudy (anonymous):



OpenStudy (anonymous):

using pythagorean theorem

OpenStudy (anonymous):

would it be a^2+b^2+c2

OpenStudy (anonymous):

Sorry, a^2+b^2=c^2

OpenStudy (anonymous):

7^2+25^2=c^2 49+625=c^2 c^2=674 c=25.96

OpenStudy (anonymous):

@killerdime20

OpenStudy (anonymous):

no the 25 is c

OpenStudy (anonymous):

a and b are the legs and c is the hypotenuse

OpenStudy (anonymous):